# DeepTalk: Towards Seamless and Smart Speech Interaction with Adaptive Modality-Specific MoE

### v2

@@ C.3 Modality-Specific MoE
Model LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S PureMoE 48.3 40.7 20.9 15.7 21.5 17.6 30.2 24.7 Random Modality-Specific(1) 40.3 32.7 17.3 10.9 18.9 14.5 25.5 19.4 Random Modality-Specific(2) 54.3 45.7 25.9 19.8 26.3 20.1 35.5 28.5 Random Modality-Specific(3) 52.1 47.0 23.3 17.6 29.8 23.4 35.1 29.3 Adaptive Modality-Specific 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
As shown in Table 8, we compared two settings: PureMoE without specifying modality experts and Random Modality-Specific MoE with randomly assigned experts. We found that if modality experts are not set, the model can only rely on self-learning the modality boundaries. During the early stages of training, the model does not clearly understand these boundaries, and the inclusion of audio data may significantly harm important text experts, leading to poorer final performance. The random assignment of modality experts was attempted three times, but results were inconsistent. If suitable experts were chosen as audio experts, the performance could be quite good, as seen in the second and third instances. However, if some crucial text experts were chosen as audio experts, the initial LLM suffers substantial damage, resulting in noticeable performance degradation. Using an adaptive modality assignment method effectively selects which experts are suitable as audio experts and text experts, achieving better performance.
Appendix D Border Impact
DeepTalk leverages pre-trained LLMs, which are inherently subject to the limitations of LLMs. These limitations include the potential to generate inaccurate information or biased outputs. To address these issues, we enhanced the model’s speech interaction capabilities using AudioQA and performed speech-language instruction tuning on high-quality datasets. Despite these improvements, we recommend exercising caution and conducting thorough safety and fairness evaluations before deploying the DeepTalk model in any downstream applications.

### v3

@@ B.1.3 Modality-Specific MoE
Model LLaMA Question Web Question TriviaQA Avg S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S S $\rightarrow$ T S $\rightarrow$ S PureMoE 48.3 40.7 20.9 15.7 21.5 17.6 30.2 24.7 LoRA-PureMoE 50.4 37.6 22.3 13.8 23.1 14.1 31.9 21.8 Random Modality-Specific(1) 40.3 32.7 17.3 10.9 18.9 14.5 25.5 19.4 Random Modality-Specific(2) 54.3 45.7 25.9 19.8 26.3 20.1 35.5 28.5 Random Modality-Specific(3) 52.1 47.0 23.3 17.6 29.8 23.4 35.1 29.3 Adaptive Modality-Specific 65.0 59.7 34.4 23.1 41.2 27.5 46.9 36.8
As shown in Table 8, we compared two settings: PureMoE without specifying modality experts and Random Modality-Specific MoE with randomly assigned experts. We found that if modality experts are not set, the model can only rely on self-learning the modality boundaries. During the early stages of training, the model does not clearly understand these boundaries, and the inclusion of audio data may significantly harm important text experts, leading to poorer final performance. The random assignment of modality experts was attempted three times, but results were inconsistent. If suitable experts were chosen as audio experts, the performance could be quite good, as seen in the second and third instances. However, if some crucial text experts were chosen as audio experts, the initial LLM suffers substantial damage, resulting in noticeable performance degradation. LoRA-PureMoE builds on PureMoE by applying LoRA-based fine-tuning to the LLM component. Although this mitigates the degradation of the LLM’s text capabilities, the speech modality learning remains insufficient. Using an adaptive modality assignment method effectively selects which experts are suitable as audio experts and text experts, achieving better performance.
B.2 Broader Impact
DeepOmni leverages pre-trained LLMs, which are inherently subject to the limitations of LLMs. These limitations include the potential to generate inaccurate information or biased outputs. To address these issues, we enhanced the model’s speech interaction capabilities using AudioQA and performed speech-language instruction tuning on high-quality datasets. Despite these improvements, we recommend exercising caution and conducting thorough safety and fairness evaluations before deploying the DeepOmni model in any downstream applications.
Appendix C The Use of Large Language Models
We used LLMs for grammar correction and proofreading.